# Learning from Stochastic Teacher Representations Using Student-Guided Knowledge Distillation

### v1

@@ 4.4 Discussion
Supervised vs. Unsupervised Setting. The proposed method can enhance performance both in supervised and unsupervised settings. The performance boost observed in the supervised setting is slightly higher than in the unsupervised setting; this could be because the unsupervised contrastive loss-based methods are already equipped with the ability to learn generalized representation. The augmented views of the input data in both TS2Vec and SoftCLT methods lead to generalized representations.
Why is SSD Different from just Teaching Dropout to the Student Model? Intuitively, it seems that the proposed method might be an alternative way of teaching the student model to drop out. However, during both the teacher training step and student training, the standard training-time dropout is activated, but the performance boost is not observed. The performance boost can be explained by the filtering of the teacher representations, where teaching the student to drop out would be equivalent to distilling from all stochastic representations $f^{\mathcal{T}}_{i}(x)$ instead of the attended teacher representation $\hat{f}^{\mathcal{T}}(x)$, which is then filtered through the proposed SSD method (see Appendix 0.D for a detailed discussion). This explanation is also supported by the results in Table 5 where we first distill from all $f^{\mathcal{T}}_{i}(x)$, and the student performance does not improve.
@@ 5 Conclusion
In this work, we introduced SSD, a novel approach to generating diversity in the teacher space by leveraging the stochastic nature of neural networks using distillation-time dropout and applying student-guided KD to learn meaningful representations. It employs the student’s current representation as a guide to select meaningful representations, implicitly mimicking an ensemble of task-relevant representations. Extensive experiments on real-world time-series data, complemented by validation on a benchmark vision dataset, demonstrate the effectiveness of SSD in improving representation learning. While SSD outperforms SOTA methods, its current evaluation is limited to architectures that already incorporate dropout. Given that SSD operates in the latent space, we hypothesize that the proposed student-guided KD framework could be extended to other settings where diversity is introduced through perturbations in the feature space or noise injection. This presents an exciting avenue for future research and a promising alternative to deep ensembles and explores an alternative way to learning generalized representations for time-series data, paving the way for more efficient and robust representation learning in diverse applications.
@@ 0.A.1 Datasets
CIFAR10 and CIFAR100: The CIFAR-10 and CIFAR-100 datasets are widely used benchmarks in the field of computer vision. Both datasets were created by the Canadian Institute for Advanced Research (CIFAR) and contain 60,000 32x32 color images. The CIFAR-10 dataset is composed of 10 mutually exclusive classes, each containing 6,000 images, representing a diverse set of objects such as airplanes, automobiles, birds, and cats. CIFAR-100 dataset extends this diversity by including 100 different classes, with 600 images per class, providing a more challenging problem due to the larger number of categories and increased intra-class variance.
0.A.2 Algorithm for SSD training
Algorithm 1 - SSD Training Procedure.
2:Trained student model $\mathcal{S}$.
5:for each input sample $x\in\mathcal{X}$ do
9:     Mask $\alpha_{i}$ for all indices outside the 90th percentile using Eq. (2)
11:     Compute distillation loss $\mathcal{L}_{\text{dist}}$ using Eq. (4)
12:end for
13:Compute total loss $\mathcal{L}_{\text{total}}$ using Eq. (5)
14:Update parameters $\theta^{\mathcal{S}}$ using backpropagation
0.A.3 Network Architecture for the HAR Dataset
@@
Implementation Details - HAR Dataset: For validation on Human Activity Recognition (HAR) Dataset, we use the barebones 1D CNN as shown in Table.6. The teacher model was trained for a total of 100 epochs with Adam optimizer. The initial learning rate was 0.05 with a ReduceLROnPlateau scheduler on training loss with patience of 10 and a reduction factor of 0.1. While training the student, the total number of repetitions to generate teacher representations was 30. The batch size was 128, and the distillation-time dropout rate was 0.2.
0.A.4 Validating SSD Beyond Time-Series
@@
Appendix 0.B Train Time Computation Comparison
@@ 0.C.1 Effect of Student Parameters Initialization with Teacher Weights
SSD 0.9182
The current student representation is what guides the distillation process at each distillation step. This section investigates the impact to student parameter initialization with the trained teacher weights $\theta^{\mathcal{T^{\prime}}}$. Tab. 8 compares the performance of SSD on the HAR dataset. The baseline i.e. the teacher model achieves $0.9002$ accuracy score. However the student model performance with random initialization is 0.8975, which is lower than the baseline performance. As explained in section 4, this decrease in performance is observed because the distillation process is guided by the current student representation $f^{\mathcal{S}}$. If the student model is not initialized with the teacher weights, it learns a separate trajectory to converge. This phenomenon is similar to the concept of pretraining in weight averaging methods, where if weights of different models trained from scratch are averaged, the performance diminishes drastically because each model learns a different trajectory towards convergence. On the other hand, if the models start with pre-trained weights, they are more likely to converge with similar weights, and averaging them leads to better performance. Similarly, in SSD, if the student model parameters are randomly initialized, the student representation $f^{\mathcal{S}}$ for each $x\in X$ will be misaligned with the teacher representation. It is only after the student model parameters with $\theta^{\mathcal{T}}$, that the $\phi_{i}$ calculated is aligned leading to a teacher representation $\hat{f}^{\mathcal{T}}$ that has meaningful diversity aligned with the student representation.
@@
Tab. 9 also presents similar results where the average of the 12 selected datasets from the UCR archive, drops from 0.8426 to 0.8125 with random initialization of student parameters. Whereas the student model with teacher weights initialization achieves 0.8508 average accuracy score.

### v2

@@ 4.4 Discussion
Supervised vs. Unsupervised Setting. The proposed method can enhance performance both in supervised and unsupervised settings. The performance boost observed in the supervised setting is slightly higher than in the unsupervised setting; this could be because the unsupervised contrastive loss-based methods are already equipped with the ability to learn generalized representation. The augmented views of the input data in both TS2Vec and SoftCLT methods lead to more generalized representations.
Why is SSD Different from just Teaching Dropout to the Student Model? Intuitively, it seems that the proposed method might be an alternative way of teaching the student model to drop out. However, during both the teacher training step and student training, the standard training-time dropout is activated, but the performance boost is not observed. The performance boost can be explained by the filtering of the teacher representations, where teaching the student to drop out would be equivalent to distilling from all stochastic representations $f^{\mathcal{T}}_{i}(x)$ instead of the attended teacher representation $\hat{f}^{\mathcal{T}}(x)$, which is then filtered through the proposed SSD method (see Appendix D for a detailed discussion). This explanation is also supported by the results in Table 5 where we first distill from all $f^{\mathcal{T}}_{i}(x)$, and the student performance does not improve.
@@ 5 Conclusion
In this work, we introduced SSD, a novel approach to enhancing diversity in the teacher space by leveraging the stochastic nature of DL models using distillation-time dropout and applying SGKD to learn meaningful representations. It employs the student’s current representation as a guide to select meaningful representations, implicitly mimicking an ensemble of task-relevant representations. Extensive experiments on real-world time-series data, complemented by validation on a benchmark vision dataset, show the effectiveness of SSD in improving representation learning. While SSD outperforms SOTA methods, its current evaluation is limited to architectures that already incorporate dropout. Given that SSD operates in the latent space, we hypothesize that the proposed SGKD framework could be extended to other settings where diversity is introduced through perturbations in the feature space or noise injection. This presents an compelling avenue for future research and a promising alternative to deep ensembles. It also provides an alternative for learning generalized representations for time-series data, paving the way for more efficient and robust representation learning in a wide range of applications.
{credits}
5.0.1 Acknowledgements
This research endeavor was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), Fonds de recherche du Québec – Santé (FRQS), Canada Foundation for Innovation (CFI), and the Digital Research Alliance of Canada.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
@@ 0.A.1 Datasets
CIFAR10 and CIFAR100: The CIFAR-10 and CIFAR-100 datasets are widely used benchmarks in the field of computer vision. Both datasets were created by the Canadian Institute for Advanced Research (CIFAR) and contain 60,000 32$\times$32 color images. The CIFAR-10 dataset is composed of 10 mutually exclusive classes, each containing 6,000 images, representing a diverse set of objects such as airplanes, automobiles, birds, and cats. CIFAR-100 dataset extends this diversity by including 100 different classes, with 600 images per class, providing a more challenging problem due to the larger number of categories and increased intra-class variance.
0.A.2 Network Architecture for the HAR Dataset
@@
Implementation Details - HAR Dataset: For validation on Human Activity Recognition (HAR) Dataset, we use 1D CNN barebones as shown in Table 6. The teacher model was trained for a total of 100 epochs with Adam optimizer. The initial learning rate was 0.05 with a ReduceLROnPlateau scheduler on training loss with patience of 10 and a reduction factor of 0.1. While training the student, the total number of repetitions to generate teacher representations was 30. The batch size was 128, and the distillation-time dropout rate was 0.2.
0.A.3 Validating SSD Beyond Time-Series
@@
Appendix 0.B Comparison of Train Time Complexity
@@ 0.C.1 Effect of Student Parameters Initialization with Teacher Weights
SSD (ours) 0.9182
The current student representation is what guides the distillation process at each distillation step. This section investigates the impact to student parameter initialization with the trained teacher weights $\theta^{\mathcal{T^{\prime}}}$. Table 8 compares the performance of SSD on the HAR dataset. The baseline i.e. the teacher model achieves $0.9002$ accuracy score. However the student model performance with random initialization is 0.8975, which is lower than the baseline performance. As explained in Section 4.3, this decrease in performance is observed because the distillation process is guided by the current student representation $f^{\mathcal{S}}$. If the student model is not initialized with the teacher weights, it learns a separate trajectory to converge. This phenomenon is similar to the concept of pretraining in weight averaging methods, where if weights of different models trained from scratch are averaged, the performance diminishes drastically because each model learns a different trajectory towards convergence. On the other hand, if the models start with pre-trained weights, they are more likely to converge with similar weights, and averaging them leads to better performance. Similarly, in SSD, if the student model parameters are randomly initialized, the student representation $f^{\mathcal{S}}$ for each $x\in X$ will be misaligned with the teacher representation. It is only after the student model parameters with $\theta^{\mathcal{T}}$, that the $\phi_{i}$ calculated is aligned leading to a teacher representation $\hat{f}^{\mathcal{T}}$ that has meaningful diversity aligned with the student representation.
@@
Table 9 also presents similar results where the average of the 12 selected datasets from the UCR archive, drops from 0.8426 to 0.8125 with random initialization of student parameters. Whereas the student model with teacher weights initialization achieves 0.8508 average accuracy score.
0.C.2 Effect of Percentile Thresholding on Performance
The proposed SSD method relies on the dynamic selection of the teacher representation based on percentile thresholding. The parameter $\epsilon$ controls the percentile value. This section studies the effect of different $\epsilon$ values on the overall student performance.
Table 10: Effect of $\epsilon$ on the student performance
$\epsilon$ 20 30 50 60 70 80 90 100
Accuracy 0.9024 0.9082 0.9082 0.9085 0.9152 0.9152 0.9182 0.9148
Table 10 shows the student performance with increasing value of $\epsilon$. For lower values of $\epsilon$, the student performance is close to the baseline. As $\epsilon$ increases, the far-off representations in the teacher space are effectively masked/filtered, allowing only relevant ones to be included in the $\hat{\alpha}_{i}$. The model’s performance peaks at $\epsilon=90$. A decrease in performance is observed again at $\epsilon=100$, likely because this high threshold selects only the single teacher representation most similar to the student representation, significantly limiting diversity in the teacher and the student representations. Another interesting observation is that the performance is not as sensitive to the $\epsilon$ value as it is to statically selecting top-$k$ representations. This is because the top-$k$ is unaware of batch dynamics always selects a $k$ number of representations, whereas $\epsilon$-thresholding dynamically selects only the relevant teacher representations.